OpenStudy (anonymous):

Evaluate the expression 2a + 4b – 3c for a = 9, b = 5, c = 12Select one of the options below as your answer: A. 5 B. 84 C. 2 D. 1,284

OpenStudy (anonymous):

=2(9)+4(5)-3(12) =18+20-36 =38-36 =2

OpenStudy (anonymous):

thx